Week of 161205

WLCG Operations Call details

  • At CERN the meeting room is 513 R-068.

  • For remote participation we use the Vidyo system. Instructions can be found here.

General Information

  • The purpose of the meeting is:
    • to report significant operational issues (i.e. issues which can or did degrade experiment or site operations) which are ongoing or were resolved after the previous meeting;
    • to announce or schedule interventions at Tier-1 sites;
    • to inform about recent or upcoming changes in the experiment activities or systems having a visible impact on sites;
    • to provide important news about the middleware;
    • to communicate any other information considered interesting for WLCG operations.
  • The meeting should run from 15:00 until 15:20, exceptionally to 15:30.
  • The SCOD rota for the next few weeks is at ScodRota
  • General information about the WLCG Service can be accessed from the Operations Web
  • Whenever a particular topic needs to be discussed at the daily meeting requiring information from site or experiments, it is highly recommended to announce it by email to wlcg-operations@cernSPAMNOTNOSPAMPLEASE.ch to make sure that the relevant parties have the time to collect the required information or invite the right people at the meeting.

Tier-1 downtimes

Experiments may experience problems if two or more of their Tier-1 sites are inaccessible at the same time. Therefore Tier-1 sites should do their best to avoid scheduling a downtime classified as "outage" in a time slot overlapping with an "outage" downtime already declared by another Tier-1 site supporting the same VO(s). The following procedure is recommended:

  1. A Tier-1 should check the downtimes calendar to see if another Tier-1 has already an "outage" downtime in the desired time slot.
  2. If there is a conflict, another time slot should be chosen.
  3. In case stronger constraints cannot allow to choose another time slot, the Tier-1 will point out the existence of the conflict to the SCOD mailing list and at the next WLCG operations call, to discuss it with the representatives of the experiments involved and the other Tier-1.

As an additional precaution, the SCOD will check the downtimes calendar for Tier-1 "outage" downtime conflicts at least once during his/her shift, for the current and the following two weeks; in case a conflict is found, it will be discussed at the next operations call, or offline if at least one relevant experiment or site contact is absent.

Links to Tier-1 downtimes

ALICE ATLAS CMS LHCB
  BNL FNAL  

Monday

Attendance:

  • local: Xavier (IT-ST/SCOD), Vincent (Security), Marian (WLCG), Emil (DB), Sabine (ATLAS), Maarten (ALICE), Julia, Maria (GGUS), Stephan (CMS)
  • remote: Luca (INFN), Andrew (NL-T1), John (RAL), FaHui (ASGC), Antonio(LHCb), Christian (NDGF), Di Qing (Triumf), David (FNAL), Stephan (CMS), Di (TRIUMF), Dimitri (KIT), David (FNAL), Rolf (IN2P3), Pepe (PIC), Stefano (CMS), Kyle (OSG), Xin (BNL)

Experiments round table:

  • ATLAS reports ( raw view) -
    • Production
      • Big production of DAOD derivations on 2015/2016 pp data and MC and urgent upgrade task ongoing. Some difficulties to keep the share requested since upgrade task need high memory queue and derivation tasks take time to merge files.
      • Some 2016 data reprocessing tests with release 21 starting
      • High level of jobs running : 200-260k running job slots over the week
      • One misconfigured Powheg task that should use single core but was spawning over all the cores: problem known and fixed that reappeared because jobs was extension of an old dataset
    • Services
      • Migration of virtual machines ongoing (week 05 December - 09 December)
      • still CVMFS probe (stratum 1 and comb db) going up and down (behind revision problem)

  • CMS reports ( raw view) -
    • Big production campaign for 2017 winter conferences in progress. Using all resources.
    • Some bottlenecks removed from user analysis submissions. Users submitting record amount to work but squeezed out of resources by production.
    • Staging from tape is still a clear limitation, in some cases we have found it faster to re-create samples, but of course there is a waste. Discussion with sites in progress, admins are helpful and rates have improved at KIT and JINR. But we are not out of the woods yet and we may have to review procedures on both our and their side for the future.
    • All in all : signs of a healthy system being uses at capacity and stressed to the limits, which is where we want to be. Of course we want to push limits further away.
    • No new major problem.

  • ALICE -
    • Successful conclusion of this year's data taking.
      Thanks in particular to the CASTOR and EOS teams!
    • Continuing high activity, in particular for Quark Matter 2017, Feb 5-11

  • LHCb reports ( raw view) -
    • Activity
      • Production activity running fine, nothing to report.
    • Site Issues
      • T0: NTR
      • T1:
        • Online: Some transfer issues.

Sites / Services round table:

  • ASGC: ntr
  • BNL: NTR
  • CNAF: NTR
  • EGI: NTR
  • FNAL: NTR
  • GridPP:
  • IN2P3: Last reminder: downtime tomorrow, as already announced before
  • JINR:
    • Backlog in tape reading queue dropped to 4000 in PhEDEx stage-in agent. Last day transfer rate was around 270MB/sec.
    • PhEDEx upgrade planned later in December, we will wait for the exhaustion of tape reading queue.
  • KISTI:-
  • KIT: NTR
  • NDGF: No IPv6 access to lcg-bdii.cern.ch? Downtime Thursday for headnode update. Should be invisible due to redundant setup.
  • NL-T1: on Nov 27th around 1800 UTC, the NIKHEF-ELPROD section of the facility suffered filesystem corruption of the transient jobs areas. Reconstruction having been ineffective, all jobs had to be flushed and pending local (stdout/stderr/sandboxes) were lost. Service was restored ~4hrs later on 22:37 UTC. Apologies for the jobs lost. Network testing SARA 15th December
  • NRC-KI:
  • OSG: Received the tickets (total of 12) from CMS requiring to upgrade Phedex instances.
  • PIC: NTR
  • RAL: There was a PDU failure this morning (05/12/2016 ~1:45) which caused some of the RAL infrastructure to lose connectivity. Staff attended site and restored services. There were a total of 3 disk servers (one Atlas and two LHCb) unavailable at times over the weekend. One of the LHCb machines is still in intervention.
  • TRIUMF: NTR

  • CERN computing services:
  • CERN storage services: End of p-Pp run >6PB recorded in 3 weeks.
  • CERN databases: Network maintenance works on 1 link to Wigner (prolonged with the link down for over 12 hours) and resulting network congestion on the remaining link, caused all Data Guard standby replicas in Wigner to lag behind primary databases by several hours. Catching up took until 5PM for some replicas. Should had no impact for the experiments as Wigner replicas aren't really used on regular basis.
  • GGUS: The 30/11 Release went well. https://ggus.eu/?mode=didyouknow#2016-11-30 explains the list and reasons of disallowed file extensions for attachments introduced with this release. NB! Year-End info on GGUS: During the Year End holiday period, if you observe a problem with GGUS and you cannot submit a ticket via https://ggus.eu, please contact the KIT site operator (info in https://goc.egi.eu/portal). The KIT Emergency telephone number is +49 721 60828383
  • Monitoring:
    • Draft reports for the October 2016 availability sent around
  • MW Officer: NTR
  • Networks: NTR
  • Security: NTR

AOB:

Edit | Attach | Watch | Print version | History: r21 < r20 < r19 < r18 < r17 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r21 - 2016-12-05 - MaartenLitmaath
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LCG All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2019 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback